How to complete rational numbers? Mention processes.


1
Expert's answer
2021-05-07T14:11:39-0400

Rational numbers are represented in "\dfrac{p}{q}" form where q is not equal to zero. It is also a type of real number. Any fraction with non-zero denominators is a rational number. Hence, we can say that ‘0’ is

also a rational number, as we can represent it in many forms such as "\dfrac{0}{1}, \dfrac{0}{2}, \dfrac{0}{3}, \text{ etc. But, } \dfrac{1}{0}, \dfrac{2}{0}, \dfrac{3}{0}" , etc. are not rational, since they give us infinite values.


Rational numbers are expressed in the form of fractions, but we do not call them fractions as fractions include only positive numbers, while rational numbers include both positive and negative numbers.